# chattersPwCHparticipants with chatbot history

### v1

@@ 4.1.2. Straight Usage of GAI Could Lead to a Generation of Lazy Thinking Programmers
When differentiating submissions by whether they constitute a close match to generated code, our sequence analysis reveals a behavioral pattern that appears to be a clear case of work avoidance: the chatbot is prompted to generate code, some or all of the code is copied and pasted into the answer box, and a submission – with very high similarity to the generated code – is made. We found this pattern to occur most often, however, it was exhibited by less than half of all . This suggests that only some students engage in this direct form of work delegation, but those that do may try the approach multiple times.
@@
The pattern we identified from the most participants consists of an attempt with no close match to generated code, then a paste into the chatbot prompt window, and the submission of a Codegen prompt. This indicates that many participants sought code generation from the after making a submission they had at least partially thought up themselves. This is also corroborated by the preceding activity, as the majority of students made at least two submission attempts before submitting their first Codegen prompt. Although 65.2% of submitted code that constituted a close match to the ’s output at some point in the experiment, we also find that none did so on their first attempt. Our observations regarding the high prevalence of Codegen requests therefore do not show an immediate intent to “cheat” from the majority of ChatGPT-using students. Rather, it appears that only after failing to solve the problem themselves do they turn to the chatbot to delegate the remaining work.
4.2. Examples
@@
We discuss a few participant-chatbot interactions which were substantially unique.
@@ 4.2.1. GCL: The Impossible Exercise
As discussed in Section 3.2.4, the solutions generated by the GPT-3.5 chatbot were often incorrect. The chatbot responses to participant GCL were particularly misleading. Over multiple attempts, the chatbot was unable to produce the correct solution to the coding problem P1. Eventually, the chatbot claimed that the provided erroneous code sample from the exercise was actually already free of issues, and thus there was no solution to P1. This led the participant to conclude that the goal of the study was to present them with an unsolvable task [GCL-M33], that they had been tricked [GCL-M37], and that there was a connection to the field of psychology [GCL-M43]. At this point, the participant closed the session and ended the experiment.
This interaction shows that a conversational generative AI agent may convince a learner that the exercise itself is faulty, rather than that the agent is incapable of generating a correct solution – despite the prominence of warnings in tools like ChatGPT not to trust their output and validate claims through external sources. Whether students are generally aware of these limitations, or consider ChatGPT to be speaking with authority, is not clear.
While only participant GCL was directly told by the chatbot that the exercise was impossible, other participants also received responses that contradicted the output of the test runner, such as reassuring a student that the generated solution would work even though it hadn’t [E6R-M15]. It remains unclear whether other participants concluded the experiment under the impression that the task itself was faulty or unsolvable, or as they thought the were unable to solve the problem.
@@
This exchange highlights a problem we briefly discussed above: generated code can incorporate concepts that a novice learner may not be familiar with yet. In our analysis of ChatGPT’s ability to pass the course, this was only relevant for the student’s ability to explain their work during code review. Here, it had a much more immediate impact, as the student had to navigate a novel concept entirely on their own, potentially not even realizing that they were missing crucial context. The student thus struggled to reconcile an error message they could not understand with a chatbot claiming that the error was fixed, without any guidance on what their mistake was.
4.3. Implications and Future Work
Our findings suggest that once students decide turn to a ChatGPT-like assistant for help, especially if they already have prior experience with generative AI, they predominantly seek code solutions rather than support in creatively working through the problem (Wei et al., 2024). As we did not test the students’ understanding of their final submission after completion, we cannot ascertain what students retained from the exercise, especially for those who were largely guided by the chatbot. Whether learners can actually benefit from with their self-regulated usage strategies therefore warrants further research.
Additionally, we believe the perception of and ChatGPT in particular should be investigated further, to understand what motivates the students’ behavior and impact their prompting skills (Kruse et al., 2024). In our study, we created an environment that was as close to the students’ regular working environment as possible, to elicit their typical behavior. This means we did not capture a lot of information on the students’ thoughts and impressions during the experiment. Future studies could use exit surveys or think-aloud observations to gain more insight into student sentiment while using tools for programming (Kruse et al., 2024).
Our ChatGPT-like chatbot did not include any task-specific instructions in the system prompt to elicit certain behavior from the model. Different behaviors could encourage students to think through the problem themselves or “nudge” them towards finding the solution (Pham et al., 2021), potentially even beyond the model’s inherent reasoning capabilities. A follow-up study could present students with a problem that the is unable to generate correct code for, and compare different prompt instructions by whether they can still improve the students’ own task-solving performance.
4.4. Limitations
4.4.1. Task-Solving Evaluation
@@
The temperature parameter of the was set to 0.0 to enhance replicability. However, as the models are provided through an external service, they may be updated and yield different outputs for the configuration and prompts used in this work. Some studies have reported significant changes in the behavior of GPT-3.5 and GPT-4 over time (Chen et al., 2023), though it is not clear whether these results are based on architectural changes, continuous training, or simply the non-deterministic nature of the .
@@
Many of the exercises in the studied course are adaptations of those found in Barnes and Kölling (Barnes and Kölling, 2009), an often cited book for teaching object-oriented programming using Java and BlueJ. We expect the performance results on the writing and in-person coding exercises to be generalizable to other courses that have drawn from the same sources when designing their exercises, aside from the atypical code style conventions of the studied course. However, the phrasing of the prompt and problem statement can significantly impact whether the will solve it correctly (Denny et al., 2023). The logical correctness of model responses may therefore be reflective of the way the problem statements in the course are written and not the ’s actual ceiling for reasoning.
Furthermore, the material used is entirely in German, which may affect the performance of the language models (Lai et al., 2023). To gauge whether this would have a significant impact on our results, we reviewed some of the problems the were unable to solve and manually prompted them with English translations. This yielded responses of comparable quality, with the same logical errors.
4.4.2. Experiment
@@
Participants may have been influenced by the experimental setting or the knowledge of their actions being recorded when deciding whether to use the chatbot. The same potential risk exists for the self-reported use of generative AI, which students may have felt discouraged from sharing honestly. They may have thought it to be more socially acceptable to report slightly lower than true ChatGPT usage numbers due to the perception of such activity as engaging in cheating (Zastudil et al., 2023). However, as the results are fairly consistent with similar surveys in recent months (Kazemitabaar et al., 2023; Prather et al., 2023a; Zastudil et al., 2023) and we were commonly able to observe behavior that could be considered unethical in an academic context (Prather et al., 2023a), we do not believe this to be a major limitation.
@@
The participant pool was comprised entirely of students of one university course in one year. Sentiments and strategies on use may vary across different populations and over time, limiting the generalizability of our findings. Thus, replicating our study in different universities is desirable.
Finally, only a single two-part exercise was used to analyze the students’ behavior. The exercise was rather complex requiring comprehension and thinking. The propensity of learners to use outputs in an exercise has been shown to vary significantly by exercise topic (Kazemitabaar et al., 2023). Therefore, the same study design may yield different results with a simpler algorithmic code authoring task.
@@ 5.1. Overall Sentiment
Kasneci et al. (Kasneci et al., 2023) summarize various challenges posed by the use of in education. The authors particularly highlight the risk of learners relying too heavily on the model, noting that “the model simplifies the acquisition of answers or information, which can amplify laziness and counteract the learners’ interest to […] come to their own conclusions and solutions” (Kasneci et al., 2023, p. 5). Our work is the first to provide quantitative empirical evidence and detailed behavioral insights detailing such concerns specifically for programming education.
Lau and Guo (Lau and Guo, 2023) interviewed instructors of university introductory programming courses on their sentiment towards the use of ChatGPT in education. They found that most educators were very unsure how many of their students actually used ChatGPT. Notably, every single interviewee of their study independently brought up cheating concerns near the start of their interview, indicating that this is an important issue for educators in the space. These concerns have been echoed by a smaller interview study by Zastudil et al. (Zastudil et al., 2023), where all instructors and a majority of the students reported anticipating an increase in academic dishonesty due to tools. Similar sentiment has also been shown in a survey by Prather et al.  (Prather et al., 2023a). They found 51% of surveyed instructors to believe either many or almost all students “are using GenAI Tools in ways that [they] would not approve of” (Prather et al., 2023a, p.13). Almost every educator considered submission of generated code unethical if the student did not understand it, though only 60% would have also considered it unethical if the student had taken the time to read and understand the auto-generated solution first (Prather et al., 2023a).
In response, some instructors have begun shifting their course grade composition more towards exam scores, explicitly showing students the limitations of AI code generation, or outright banning ChatGPT and comparable tools in their class (Lau and Guo, 2023, p. 113). Zastudil et al. saw students universally disagreeing with banning tools outright (Zastudil et al., 2023), though Prather et al. found that the majority of students support at least some level of restrictions on its use (Prather et al., 2023a).
@@ 5.2. Task-Solving Performance
Finnie-Ansley et al. (Finnie-Ansley et al., 2022, 2023) analyzed the performance of the Codex model (Chen et al., 2021) on educational programming problems of various difficulties. For questions on a CS1 course assessment, they found that Codex solved 43.5% of questions correctly on the first try, ignoring trivial formatting errors; when given multiple attempts with the same penalty scheme applied to students, the Codex model would have placed in the top quartile of students (Finnie-Ansley et al., 2022). In a subsequent analysis of exam questions in a CS2 course, the researchers also found Codex placing in the top 25% of the students (Finnie-Ansley et al., 2023). Popovici (Popovici, 2023) found ChatGPT able to solve 68% of coding exercises of a functional programming course in the first try, expanding to 86% after follow-ups. They found similar performance for “easy” and “medium” difficulty problems, but fewer correct answers for “hard” problems.
Berrezueta-Guzman and Krusche (Berrezueta-Guzman and Krusche, 2023) analyzed ChatGPT’s performance on the programming tasks of their introductory university course. They found that a student could pass the course using only ChatGPT-generated answers, albeit with a grade of 55%, only scoring slightly above the passing threshold. They also analyzed the time required to copy-paste the question text into the ChatGPT input window as well as inserting the response into the answer field. They found a 91% reduction in the time invested in solving the assignments naively, i.e. simply submitting the first generated solution, and still a 74% reduction in time when the code was then manually adjusted to receive full marks (Berrezueta-Guzman and Krusche, 2023, p. 149).
@@ 5.3. Interactions of Programming Students and Programmers with
MacNeil et al. (MacNeil et al., 2022, 2023) demonstrated a potential application for in generating beginner-friendly explanations for source code. Overall, the explanations were mostly rated useful by students, both for their own understanding and as a general learning tool (MacNeil et al., 2023). We observed that without prior guidance and instructions, students barely used the bot for explanation. Leinonen et al. (Leinonen et al., 2023) found that automatically generated code explanations by GPT-3 were rated as both a more accurate and easier to understand than explanations created by students. However, unlike MacNeil et al., they found students to prefer line-by-line explanations (Leinonen et al., 2023).
@@
Kazemitabaar et al. (Kazemitabaar et al., 2023) studied the behavior of learners when given access to the Codex model, and the impact of the tool on their learning progression. For questions related to loops, 60% of submitted answers in the Codex group were entirely AI-generated (Kazemitabaar et al., 2023, p. 10). On follow-up tasks without access to a code generator, the authors found no significant difference in the correctness of solutions presented by learners who had and had not used Codex earlier. In an assessment a week later, learners in the Codex group exhibited significantly higher rates of errors on code authoring tasks with no starter code given, but had similar correctness scores and completion times (Kazemitabaar et al., 2023).
Prather et al. (Prather et al., 2023b) studied the interactions of novice programming students with GitHub Copilot, an autocompletion-style code generation tool. They observed a common pattern of “drifting”, where participants would be led astray by incorrect generated solutions which they would have to correct, as well as “sheperding”, where they would focus more of their attention on guiding the towards suggesting a correct solution than writing it on their own (Prather et al., 2023b).
@@
Hellas et al. (Hellas et al., 2023) prompted ChatGPT and Codex with real help requests from students of an online programming course and assessed the models’ ability to identify the problem and provide meaningful assistance. While ChatGPT-3.5 was able to identify all issues in the majority of code snippets, it notably also found non-existent issues in 40% of cases.
@@
While much research has been conducted on use cases, performance evaluation, and limitations of for code generation and software development in general, studies of how developers interact with (including behavioral and prompting patterns) are still sparse. Most notable, Jin et al. (Jin et al., 2024) recently analyzed prompting patterns and code reuse in the DevGPT dataset (Xiao et al., 2024) of ChatGPT-conversations referenced on GitHub. They found generated code to be present in subsequent commits from 17% of conversations, and a further 26% with some modifications. In a user study with developers, Vaithilingam et al.  (Vaithilingam et al., 2022) found GitHub Copilot to have little impact on task completion time, but strong user preference for Copilot over the IDE’s built-in suggestions. Understanding, evaluating and debugging generated code was found to be a major challenge for participants. Analyzing ChatGPT-generated answers to StackOverflow questions, Kabir et al. (Kabir et al., 2024) found just over half of the generated answers to contain incorrect information, which was then overlooked by human evaluators in some cases. Most recently, Kruse et al. (Kruse et al., 2024) conducted an experiment on prompting skills comparing experienced programmers with students. The authors observed that prompts, particularly by students, led to lower quality generated documentation than a prepared prompt executed with a single click. Similar to our finding, the authors concluded that prompting skills cannot be expected for effective use of tools and need to be taught similarly as teaching how to use a debugger or a test suite.

### v2

@@ 4.1.2. Learning to Think with and to Develop, Compare, and Challenge Solution Paths
The most common behavioral pattern we identified in our sequence analysis appears to be a case of work avoidance: the chatbot is prompted to generate code, some or all of the code is copied and pasted into the answer box, and a submission—with very high similarity to the generated code—is made. While this pattern occurred most often, it was exhibited by less than half of all . This suggests that only some students engage in this direct form of work delegation, but those that do may try the approach multiple times. More generally, we found that the participants changed the code significantly more between submissions if they requested code generation immediately prior. In the majority of those cases, the result was more similar to the generated code than to the previous submission. This indicates that students have a tendency to align their code with output, even if they have to discard much of their existing attempt. While it is generally good to explore various thoughts and solution approaches, it is also important to be able to focus and forward-reason about an emerging solution path. Methodological as well as tool-assisted approaches to develop and sharpen such skills are yet to be explored.
@@
The pattern we identified from most participants consists of an attempt with no close match to generated code, then a paste into the chatbot prompt window, and the submission of a Codegen prompt. This indicates that many participants sought code generation from the after making a submission they had at least partially thought up themselves. This is also corroborated by the preceding activity, as the majority of students made at least two submission attempts before submitting their first Codegen prompt. Although 65.2% of submitted code that constituted a close match to the ’s output at some point in the experiment, we also find that none did so on their first attempt. Our observations regarding the high prevalence of Codegen requests therefore do not show upfront intent to “cheat” from the majority of ChatGPT-using students. Rather, it appears that only after failing to solve the problem themselves—a situation developers encounter frequenter in their daily work—do the students turn to the chatbot to delegate the remaining work.
4.1.3. On Usefulness and Usability of Tools for Novices
While reviewing the students’ interactions, we noted that multiple participants ended up in cycles of seeing an error message, prompting the chatbot for assistance, implementing its suggestion, and receiving another error message. For instance, the interaction timeline in Figure 9 shows an accelerating back-and-forth between prompts and submissions. Through our sequence analysis, we were able to quantify this interaction pattern, which turned out to be the most common across all . As the GPT-3.5 model was unable to produce a correct solution to problem P1 in most cases, even in response to Fix prompts, this was a notably ineffective strategy. Nonetheless, some participants repeatedly tried ”coercing” the model into generating a correct submission through a series of Codegen prompts. This suggests they were unable to recognize when a tool is incapable of providing effective assistance, similar to observations by Prather et al. (2024) on novices using inline code completion.
Using for code generation shifts the developer’s focus from ideation (Wei et al., 2024) to parsing and debugging unfamiliar code. Prior work has shown that this is a challenge for professional developers (Sarkar et al., 2022; Vaithilingam et al., 2022). Without the technical experience, it is even more difficult for beginners to review ChatGPT’s output. We expect this effect to be exacerbated if students also depend on ChatGPT for knowledge and conceptual questions, which has been found to be negatively associated with performance in recent work (Mailach et al., 2025). Further, tools can increase developers’ perceived productivity (Ziegler et al., 2024) and level of understanding (Prather et al., 2024) without delivering actual improvements, potentially leading to a false sense of confidence (Lee et al., 2025) among novices. Research has yet to explore effective feedback strategies that make model uncertainty transparent to users and minimize overconfidence.
We observed that ChatGPT often apologized and occasionally even justified its mistakes before attempting to correct itself, which may have increased the students’ trust in the system. This happened in 10 experiment sessions, in fact up to 14 times in one single session (participant 28X). Explanations of potential errors in an automated system have been shown to increase trust and reliance on the system, largely mitigating the negative effect of observing an error (Dzindolet et al., 2003). Thus, GPT’s acknowledgment of mistakes may actually make students less likely to break out of the error-prompting cycle. This highlights the need to sensitize students to the risks and limitations of using . Our experiment task could be used as a practical example in class, allowing students to experience first-hand how consistently ChatGPT can make mistakes, and how confidently it can appear to defend them.
4.2. Examples of Student-Bot Interactions
@@
We highlight a few student-chatbot interactions which were substantially unique.
@@ 4.2.1. GCL: The Impossible Exercise
As discussed in Section 2.2.2, the experiment task was designed to cause GPT-3.5 to produce incorrect outputs. The chatbot responses to participant GCL were particularly misleading. Over multiple attempts, the chatbot was unable to produce the correct solution to the coding problem P1. Eventually, the chatbot claimed that the provided erroneous code sample from the exercise was actually already free of issues, and thus there was no solution to P1. This led the participant to conclude that the goal of the study was to present an unsolvable task [GCL-M33], that they had been tricked [GCL-M37], and that there was a connection to the field of psychology [GCL-M43]. At this point, the participant closed the session and ended the experiment.
This interaction demonstrates a unique new challenge for learners: a conversational agent may convince them that the task itself is flawed, rather than recognizing that the agent is incapable of generating a correct solution. While only participant GCL was directly told by the chatbot that the exercise was impossible, other participants also received responses that contradicted the output of the test runner, such as reassuring a student that the generated solution would work even though it hadn’t [E6R-M15]. We do not know whether other participants concluded the experiment under the impression that the task itself was faulty or unsolvable, or that they were unable to solve the problem. Though this was an isolated incident, it raises the concerning possibility that an inexperienced student—already struggling to assert the quality of ChatGPT’s output—may be encouraged by the chatbot to trust it over their instructor on conflicting statements. While tools like ChatGPT include warnings not to trust their generations and validate claims through external sources, whether students take them seriously or consider ChatGPT to be speaking with authority warrants further investigation.
@@
In copy-pasting the generated code, the student inadvertently used an unfamiliar concept, potentially without even realizing they were missing crucial context. This is the same problem we observed in the Learn category of our task-solving evaluation, demonstrating that it can negatively impact beginners beyond the context of interview assessments. Here, the student had to reconcile an error message they could not understand with a chatbot claiming that the error had been resolved, without any guidance on what their mistake was. It is unclear how much awareness of limitations would have helped this particular student, as the generated output was correct and iterating over it further was counterproductive.
4.3. Limitations and Recommendations for Improvement
4.3.1. Threats to Validity
@@
In our task-solving evaluation, the referenced may be updated and yield different outputs for the configuration and prompts used in this work. The temperature parameter of the models was set to 0.0 to mitigate this as much as possible. However, some studies have reported significant changes in the behavior of GPT-3.5 and GPT-4 over time (Chen et al., 2023), though it is not clear whether these results are based on architectural changes, continuous training, or simply the non-deterministic nature of the .
@@
Many of the exercises in the studied course are adaptations of those found in Barnes and Kölling (Barnes and Kölling, 2009), an often cited book for teaching object-oriented programming using Java and BlueJ. We expect the performance results on the writing and in-person coding exercises to be generalizable to other courses that have drawn from the same sources when designing their exercises, aside from the atypical code style conventions of the studied course. The phrasing of the prompt and problem statement can significantly impact whether the will solve it correctly (Denny et al., 2023). The logical correctness of model responses may therefore be reflective of the way the problem statements in the course are written and not the ’s actual ceiling for reasoning. Furthermore, the material used is entirely in German, which may affect the performance of the language models (Lai et al., 2023). To gauge whether this has a significant impact on our results, we reviewed some of the problems the were unable to solve and manually prompted them with English translations. This yielded responses of comparable quality with the same logical errors.
@@
In the experiment, participants may have been influenced by the setting or the knowledge of their actions being recorded when deciding whether to use the chatbot. The same potential risk exists for the self-reported use of generative AI, which students may have felt discouraged from sharing honestly. They may have thought it to be more socially acceptable to report slightly lower than true ChatGPT usage numbers due to the perception of such activity as engaging in cheating (Zastudil et al., 2023). However, as the results are fairly consistent with similar surveys around the time (Kazemitabaar et al., 2023; Prather et al., 2023a; Zastudil et al., 2023) and we were commonly able to observe behavior that could be considered unethical in an academic context (Prather et al., 2023a), we do not believe this to be a major limitation.
@@
The participant pool consisted entirely of students of one university course in one year. Sentiments and strategies on use may vary across different populations and over time, limiting the generalizability of our findings. Thus, replicating our study in different universities is desirable.
4.3.2. Recommendations for Future Studies
Our findings suggest that once learners decide turn to a ChatGPT-like assistant for help, especially if they already have prior experience with generative AI, they predominantly seek code solutions rather than support in creatively working through the problem (Wei et al., 2024). In a future study, we recommend to test the students’ understanding of their final code submission after exercise completion. This can help to determine what students were able to retain from the exercise, especially if their solutions were largely guided by the chatbot.
Additionally, we believe the perception of and ChatGPT in particular should be investigated further, to understand what motivates the students’ behavior and how much authority and trust they place in these tools. In our study, we created an environment that was as close to the students’ regular working environment as possible, to elicit their typical behavior. This means we did not capture a lot of information on the students’ thoughts and impressions during the experiment. Exit surveys or think-aloud observations can be used to gain more insight into student sentiment while using tools for programming (Prather et al., 2024; Kruse et al., 2024).
Our results suggest that students are either not aware of the risks and limitations of using , or not sensitized enough to recognize them during regular use. Whether guidance on prompting, best practices and its limitations could meaningfully improve beginners’ ability to handle confidently incorrect outputs warrants investigation. A follow-up study could replicate our experiment, using the same or a comparable task, with students who have received prior training or educational material related to use. Alternatively, different chatbot behaviors could encourage students to think through the problem themselves or “nudge” them towards finding the solution (Pham et al., 2021), potentially even beyond the model’s inherent reasoning capabilities. Our ChatGPT-like chatbot did not include any task-specific instructions in the system prompt to elicit certain behavior from the model. A future study could compare the impact of different system prompts on the students’ code reuse and prompting behaviors, as well as whether the students’ own task-solving performance can be improved.
Finally, we only tested a single, fairly complex exercise, related to object states and references. The propensity of learners to use outputs has been shown to vary significantly by exercise topic (Kazemitabaar et al., 2023; Mailach et al., 2025). Therefore, we recommend replicating this study with exercises involving different concepts, with a similar relative complexity level.
@@ 5.1. Overall Sentiment
Kasneci et al. (2023) summarize various challenges posed by the use of in education. The authors particularly highlight the risk of learners relying too heavily on the model, noting that “the model simplifies the acquisition of answers or information, which can amplify laziness and counteract the learners’ interest to […] come to their own conclusions and solutions” (Kasneci et al., 2023, p. 5). Our work is the first to provide quantitative empirical evidence and detailed behavioral insights on the prevalence of chatbot overreliance specifically in programming education.
Lau and Guo (2023) interviewed instructors of university introductory programming courses on their sentiment towards the use of ChatGPT in education. They found that most educators were very unsure how many of their students actually used ChatGPT. Notably, every single interviewee of their study independently brought up cheating concerns near the start of their interview, indicating that this is an important issue for educators in the space. These concerns have been echoed by a smaller interview study by Zastudil et al. (2023), where all instructors and a majority of the students reported anticipating an increase in academic dishonesty due to tools. Similar sentiment has also been shown in a survey by Prather et al. (2023a). They found 51% of surveyed instructors to believe either many or almost all students “are using GenAI Tools in ways that [they] would not approve of” (Prather et al., 2023a, p.13). Almost every educator considered submission of generated code unethical if the student did not understand it, though only 60% would have also considered it unethical if the student had taken the time to read and understand the auto-generated solution first.
In response, some instructors have begun shifting their course grade composition more towards exam scores, explicitly showing students the limitations of AI code generation, or outright banning ChatGPT and comparable tools in their class (Lau and Guo, 2023, p. 113). Zastudil et al. (2023) saw students universally disagreeing with banning tools outright, though Prather et al. (2023a) found that the majority of students support at least some level of restrictions on its use.
@@ 5.2. Task-Solving Performance
Finnie-Ansley et al. (2022) analyzed the performance of the Codex model on educational programming problems of various difficulties. For questions on a CS1 course assessment, they found that Codex solved 43.5% of questions correctly on the first try, ignoring trivial formatting errors. When given multiple attempts with the same penalty scheme applied to students, the Codex model would have placed in the top quartile of students. In a subsequent analysis of exam questions in a CS2 course, Finnie-Ansley et al. (2023) also found Codex placing in the top 25% of students. Popovici (2023) found ChatGPT able to solve 68% of coding exercises of a functional programming course in the first try, expanding to 86% after follow-ups. They found similar performance for “easy” and “medium” difficulty problems, but fewer correct answers for “hard” problems.
Berrezueta-Guzman and Krusche (2023) analyzed ChatGPT’s performance on the programming tasks of their introductory university course. They found that a student could pass the course using only ChatGPT-generated answers, albeit with a grade of 55%, only scoring slightly above the passing threshold. They also analyzed the time required to copy-paste the question text into the ChatGPT input window as well as inserting the response into the answer field. They found a 91% reduction in the time invested in solving the assignments naively, i.e. simply submitting the first generated solution, and still a 74% reduction in time when the code was then manually adjusted to receive full marks (Berrezueta-Guzman and Krusche, 2023).
@@ 5.3. Interactions of Programming Students and Programmers with
MacNeil et al. (2022) demonstrated a potential application for in generating beginner-friendly explanations for source code. In a subsequent work, MacNeil et al. (2023) found that the explanations were mostly rated useful by students, both for their own understanding and as a general learning tool (MacNeil et al., 2023). We observed that without prior guidance and instructions, students barely used the bot for explanation. Leinonen et al. (2023) found that automatically generated code explanations by GPT-3 were rated as both a more accurate and easier to understand than explanations created by students. However, unlike MacNeil et al. (2023), they found students to prefer line-by-line explanations (Leinonen et al., 2023).
@@
Kazemitabaar et al. (2023) studied the behavior of learners when given access to the Codex model, and the impact of the tool on their learning progression. For questions related to loops, 60% of submitted answers in the Codex group were entirely AI-generated (Kazemitabaar et al., 2023, p. 10). On follow-up tasks without access to a code generator, the authors found no significant difference in the correctness of solutions presented by learners who had and had not used Codex earlier. In an assessment a week later, learners in the Codex group exhibited significantly higher rates of errors on code authoring tasks with no starter code given, but had similar correctness scores and completion times (Kazemitabaar et al., 2023).
@@
Hellas et al. (2023) prompted ChatGPT and Codex with real help requests from students of an online programming course and assessed the models’ ability to identify the problem and provide meaningful assistance. While ChatGPT-3.5 was able to identify all issues in the majority of code snippets, it notably also found non-existent issues in 40% of cases.
Prather et al. (2023b) studied the interactions of novice programming students with GitHub Copilot, an autocompletion-style code generation tool. They observed a common pattern of “drifting”, where participants would be led astray by incorrect generated solutions which they would have to correct, as well as “sheperding”, where they would focus more of their attention on guiding the towards suggesting a correct solution than writing it on their own.
Prather et al. (2024) analyzed the metacognitive challenges students encountered when using GitHub Copilot. While some students were able to solve their task faster with the tool, they found that metacognitive difficulties encountered by struggling students were not alleviated by . Instead, struggling students may even be faced with new difficulties, such as being unable to recognize when a code suggestion was helpful, or overestimating their understanding of the problem. Most recently, Mailach et al. (2025) analyzed the prompting patterns of students in a CS2 data structures course. They found solution generation to be both the most common prompt type and overall conversation intention. Similar to our observations, they found that students lacking conceptual understanding struggled with interpreting and using the chatbot output. They also found chatbot use and, more notably, code generation requests to be positively associated with solution correctness. However, this includes submissions with copy-pasted generated code, the prevalence of which is not stated.
@@
While much research has been conducted on use cases, performance evaluation, and limitations of for code generation and software development in general, studies of how developers interact with (including behavioral and prompting patterns) are still sparse. Most notable, Jin et al. (2024) recently analyzed prompting patterns and code reuse in the DevGPT dataset (Xiao et al., 2024) of ChatGPT-conversations referenced on GitHub. They found generated code to be present in subsequent commits from 17% of conversations, and a further 26% with some modifications. In a user study with developers, Vaithilingam et al. (2022) found GitHub Copilot to have little impact on task completion time, but strong user preference for Copilot over the IDE’s built-in suggestions. Understanding, evaluating and debugging generated code was found to be a major challenge for participants. Analyzing ChatGPT-generated answers to StackOverflow questions, Kabir et al. (2024) found just over half of the generated answers to contain incorrect information, which was then overlooked by human evaluators in some cases. Most recently, Kruse et al. (2024) conducted an experiment on prompting skills comparing experienced programmers with students. The authors observed that prompts, particularly by students, led to lower quality generated documentation than a prepared prompt executed with a single click. Similar to our finding, the authors concluded that prompting skills cannot be expected for effective use of tools and need to be taught similarly as teaching how to use a debugger or a test suite.